4. Legend states that Pieter Minuit, Director General of the New Netherlands colony, purchased Manhattan from Seyseys, chief of the Canarees, a Native American tribe, for the sum equivalent to $24 in 1626. If annual appreciation of 7% was earned on that investment, what would be the value of Manhattan in 2019. a. Consider SIMPLE INTEREST b. Consider COMPOUND INTEREST
Answer #1

Principal = $24

Time from 1626 to 2019 = 393 years

Rate of Appreciation = 7%

a) Using Simple Interest

SI = (P*R*T)/100

SI = (24*7*393)/100 = $660.24

Value in 2019 = 24+660.24 = $684.24

b) Using Compound interest rate

A = P(1+r/100)T

A= 24*(1.07)393

Value in 2019 will be A = $8472973966004.32